# Structure constants of Peterson Schubert calculus

### v1

@@ Abstract.
We give an explicit, positive, type-uniform formula for all (equivariant) structure constants of the Peterson Schubert calculus in arbitrary Lie types, using only information from the Cartan matrix of the corresponding root system $\Phi$. As an application, we derive a type-uniform formula for the mixed $\Phi$-Eulerian numbers.
@@ 1. Introduction
Now, let $G$ be a complex semisimple linear algebraic group with a Borel subgroup $B\subset G$. Recall that the cohomology ring $H^{*}(G/B;\mathbb{Z})$ of the flag variety $G/B$ has a distinguished basis consisting of Schubert classes [8]. Studying the product structure of these Schubert classes (sometimes also encompassing the study of analogous questions in generalized cohomology theories or equivariant settings) is the central problem in modern Schubert calculus. By Kleiman’s transversality theorem [23], the Schubert structure constants are non-negative integers. In type $A$, the Schubert polynomials in [29] are polynomial representatives of Schubert classes in the classical Borel presentation [9] of the cohomology ring of the complete flag variety. It is an outstanding problem in combinatorics and enumerative algebraic geometry to give a positive formula or a combinatorial model for the structure constants of multiplying Schubert polynomials.
There are also interesting works studying the structure constants of the cohomology ring of Peterson varieties (the so-called “Peterson Schubert calculus”). Let $\operatorname{Pet}_{G}$ denote the Peterson variety inside $G/B$ (see (2.3) for a precise definition). It admits an action of a one-dimensional sub-torus $S$ of the maximal torus $T\subset G$. Let $\sigma_{v_{I}}\in H_{S}^{*}(G/B)$ be a Schubert class indexed by some Coxeter element $v_{I}\in W$ for $I\subset\Delta$, where $W$ is the Weyl group and $\Delta$ is the set of simple roots. Let $\iota^{*}\left(\sigma_{v_{I}}\right)$ be the (equivariant) pullback of $\sigma_{v_{I}}$ along the inclusion $\iota:\operatorname{Pet}_{G}\rightarrow G/B$. While class $\iota^{*}\left(\sigma_{v_{I}}\right)$ depends on the choice of $v_{I}$, the class $p_{I}:=\frac{\iota^{*}\left(\sigma_{v_{I}}\right)}{m\left(v_{I}\right)}$ is independent of this choice after dividing by a certain intersection multiplicity $m\left(v_{I}\right)$, known as the (integral equivariant) Peterson Schubert class in the literature. In [16], Goldin, Mihalcea, and Singh show that $\left\{p_{I}\right\}_{I\subset\Delta}$ is a $c_{I,J}^{K}\in H_{S}^{*}(pt)$-basis of $H_{S}^{*}(\operatorname{Pet}_{G};\mathbb{Z})$, and that this basis is dual to the equivariant Borel–Moore homology basis constituted by the fundamental classes of Peterson Schubert cycles (see Theorem 2.5). Furthermore, their main result shows that the structure constants of multiplication $c_{I,J}^{K}\in H_{S}^{*}(pt)\cong\mathbb{Z}[t]$, with respect to the basis $\left\{p_{I}\right\}_{I\subset\Delta}$, defined by
@@
are polynomials in $t$ with non-negative coefficients.
A natural open question (asked in [20, P.43] for Lie type $A$) is the following
@@
For Lie type $A$, Goldin–Gorbutt [15] found positive combinatorial formulae for the structure constants in a successive way, while Abe–Horiguchi–Kuwata–Zeng [3] found a different combinatorial model computing these structure constants in non-equivariant cohomology based on a geometric viewpoint. Some years earlier, in type $A$, Harada–Tymoczko [20] gave a positive equivariant Monk formula for the product of a cohomology-degree-2 Peterson Schubert class (that is, $p_{\{i\}}$ for some $i\in\Delta$) with an arbitrary Peterson Schubert class, and Bayegan–Harada [6] gave an equivariant Giambelli formula to express Peterson Schubert class $p_{I}$ in terms of the ring generators $p_{\{i\}}$’s. Drellich [13] gave the equivariant Monk rule and Giambelli formula for all Lie types.
While the positivity of structure constants in Peterson Schubert calculus was recently established using the aforementioned geometric and combinatorial methods, the above Question 1.1 remains unresolved in general, and it is not known whether these constants can be calculated directly and uniformly from the intrinsic Lie-theoretic data of the associated root system. This paper addresses this problem. In this paper, we give a simple, explicit, positive, type-uniform formula for all structure constants in (1.1) of the Peterson Schubert calculus in arbitrary Lie types, which completely solves the above question in a type-uniform way. Our formula is a concise algebraic formula, determined only by the Cartan matrix. As far as we know, this is the first example of an explicit, complete, and type-uniform positive algebraic formula in general Schubert calculus. We hope that our method can shed some new light on related structure constants problems in Schubert calculus. We derive this formula from an algebraic viewpoint, which is different from the above-mentioned work. We now present our main results.
Let $C_{\Delta}:=\left(\left\langle\alpha_{i},\alpha_{j}\right\rangle\right)_{i,j\in\Delta}=(c_{ij})_{i,j\in\Delta}$ be the Cartan matrix associated with the semisimple Lie algebra $\mathfrak{g}$ of $G$. Harada–Horiguchi–Masuda [19] gave an explicit and beautiful Borel-type presentation of the equivariant cohomology ring of Peterson varieties in all Lie types; see Theorem 2.4. The following is our first result.
@@
Using the above Peterson Schubert monomials and the Harada–Horiguchi–Masuda presentation (2.6), we can expand the product of two monomials into the monomial basis representing integral equivariant Peterson Schubert classes, and compute all the structure constants in (1.1).
To state our main theorem, we first need to introduce some notations. For any $K\subset\Delta$, we use $C_{K}^{-1}]_{j,k}$ to denote the $(j,k)$-entry of the matrix $C_{K}^{-1}$, which is the inverse of the Cartan sub-matrix $C_{K}$ determined by $K$. Suppose $|\Delta|=n$, we use the notation $(\quad)_{J,K}$ to denote a $2^{n}\times 2^{n}$ matrix, with rows indexed by $J\subset\Delta$ and columns indexed by $K\subset\Delta$, and the ordering of the subsets of $\Delta$ is the lexicographical order. The following is our main theorem.
@@
With the notations as above. Suppose $I=\{i_{1},...,i_{l}\}\subset\Delta$. Then the structure constant $c_{I,J}^{K}$, defined in (1.1), is equal to $\frac{\operatorname{det}\left(C_{I}\right)\operatorname{det}\left(C_{J}\right)\left|W_{K}\right|}{\left|W_{I}\right|\left|W_{J}\right|\operatorname{det}\left(C_{K}\right)}$ multiplied to the $(J,K)$-entry of the product of matrices $(d^{K}_{i_{1},J})_{J,K}\cdots(d^{K}_{i_{l},J})_{J,K}$, where
@@
By letting $t=0$ in (1.3), we get a simple formula for all non-equivariant structure constants in Peterson Schubert calculus for $H^{*}(\operatorname{Pet}_{G};\mathbb{Z})$.
@@
Suppose $I=\{i_{1},...,i_{l}\}\subset\Delta$. Then the non-equivariant structure constant $n_{I,J}^{K}$, defined to be $c_{I,J}^{K}$ with $t=0$, is equal to $\frac{\operatorname{det}\left(C_{I}\right)\operatorname{det}\left(C_{J}\right)\left|W_{K}\right|}{\left|W_{I}\right|\left|W_{J}\right|\operatorname{det}\left(C_{K}\right)}$ multiplied to the $(J,K)$-entry of the product of matrices $(m^{K}_{i_{1},J})_{J,K}\cdots(m^{K}_{i_{l},J})_{J,K}$, where
@@
To prove the above theorems, we transform the structure constants in the second case of (1.3) into an infinite matrix series (see the proof of Theorem 3.5) related to the Coxeter adjacency matrix $A_{K}$, defined as $A_{K}:=2E-C_{K}$, where $E$ is the identity matrix. We use the classical Perron–Frobenius theorem to prove that this matrix series is convergent and related to an inverse of a matrix; see Lemma 3.7. Finally, we use the Sherman–Morrison formula to compute matrix inverse, which is simplified to the final formula; see Lemma 3.8.
It is well known that the entries of the inverse of any Cartan matrix are all non-negative (see, for example, [22, Exercise 13.8]), hence the entries of all the matrices $(d^{K}_{i,J})_{J,K}$ in Theorem 1.3 are polynomials in $t$ with non-negative coefficients. Since the entries in the product of such matrices remain polynomials in $t$ with non-negative coefficients, our formula is indeed a positive formula, which provides a new purely algebraic proof of the positivity of equivariant structure constants of Peterson Schubert calculus. We also give a simple criterion for when these structure constants are non-zero; see Corollary 3.13 and Remark 3.14.
The relationship between the (non-equivariant) Peterson Schubert calculus and mixed Eulerian numbers (which were introduced and studied by Postnikov [33]) has been actively explored very recently; see [7], [31], and [21]. These numbers compute the mixed volumes of $\Phi$-hypersimplices, see Section 4 for the precise definitions. The origins of this connection lie in the realization of the Peterson variety as a flat degeneration of a smooth projective toric variety, the $W$-permutohedral variety; see, for example, [1, 24, 25]. In [21], Horiguchi provided a combinatorial model introduced in [3] for the computation of mixed $\Phi$-Eulerian numbers when $\Phi$ is of type $A$ and derived a type-by-type computation for the mixed $\Phi$-Eulerian numbers for general Lie types by iteratively applying the (non-equivariant version of) Monk formula from Drellich in [13]. As an application of our main theorem, we derive a type-uniform formula for the mixed $\Phi$-Eulerian numbers in arbitrary Lie type.
@@
Let $\Phi$ be an irreducible root system of rank $n$. Let $c_{1},\ldots,c_{n}$ be non-negative integers with $c_{1}+\cdots+c_{n}=n$. The mixed $\Phi$}-Euler number $A_{c_{1}\cdots c_{n}}^{\Phi}$ can be computed using the following formula:
@@
Compared with the computations in [21], our formula for the mixed $\Phi$-Eulerian numbers is quite simple and direct, which avoids the need to discuss the changes in the Lie types case-by-case when considering sub-root systems as in [21, Section 7].
The remainder of the paper is organized as follows. In Section 2, we introduce the notations and some preliminary results which we use throughout this paper. In Section 3, we prove our main theorems. In Section 4, we give the application to mixed Eulerian numbers.
@@ Acknowledgments
The first author would like to express his gratitude to Hiraku Abe and Haozhi Zeng for helpful discussions. The authors thank Yibo Gao for some useful comments.
@@ 2. Preliminaries
In this section, we collect our notations and preliminary results for later use.
@@ 2.1. Flag varieties and Schubert varieties
Let $G$ be a complex connected semisimple algebraic group of rank $n$, $B\subset G$ be a Borel subgroup, and $B^{-}\subset G$ be the opposite Borel. Write $T:=B\cap B^{-}$ for the maximal torus common to both. Let $\mathfrak{g}=\operatorname{Lie}(G)$, $\mathfrak{b}=\operatorname{Lie}(B)$, and $\mathfrak{h}=\operatorname{Lie}(T)$ be the corresponding Lie algebras. Denote by $\Phi$ the root system determined by $(G,T)$, and by $\Delta=\{\alpha_{1},\dots,\alpha_{n}\}$ the simple roots specified by $(B,T)$. We abuse the notation by letting $\Delta$ also denote the index set of simple roots. The Weyl group $W:=\langle s_{i}\mid i\in\Delta\rangle$ is a finite Coxeter group. For any subset $I\subset\Delta$, we have the standard parabolic subgroup $P_{I}\subset G$, whose Weyl group is $W_{I}=\langle s_{i}\mid i\in I\rangle$, a parabolic subgroup of $W$.
@@
the Schubert and opposite Schubert varieties in $G/B$, respectively. It follows from (2.2) that the equivariant cohomology classes $\sigma_{w}\in H_{T}^{*}(G/B;\mathbb{Z})$, which are Poincaré dual to the fundamental class $\left[\Omega_{w}\right]$, form a basis $\left\{\sigma_{w}\mid w\in W\right\}$ of $H_{T}^{*}(G/B;\mathbb{Z})$ as a module over $H_{T}^{*}(pt)$, called equivariant Schubert classes.
@@ 2.2. Peterson varieties and some geometric constructions
(2.3) $\operatorname{Pet}_{G}:=\Bigl{\{}gB\in G/B\;\Big{|}\;\text{Ad}_{g^{-1}}(e)\in\mathfrak{b}\oplus{\textstyle\bigoplus_{\alpha\in\Delta}}\mathfrak{g}_{-\alpha}\Bigr{\}}.$
It is known that the Peterson variety $\operatorname{Pet}_{G}$ is always irreducible with complex dimension $n:=\operatorname{rank}G$ (see [32] or [27, Proposition 6.2]), but it is singular and even non-normal in general [26]. Peterson varieties also form a special class of a larger family of subvarieties of the flag variety, namely the Hessenberg varieties introduced in [12]; see [2] for a very nice survey.
@@
where $NV_{i}$ denotes the image of $V_{i}$ under the linear map $N\colon\mathbb{C}^{n}\to\mathbb{C}^{n}$. In this setting, Peterson varieties are also known as the regular nilpotent Hessenberg varieties with Hessenberg function $h(i)=i+1$ for $i=1,2,\ldots,n-1$.
It is known that there is a one dimensional sub-torus $S\subset T$ acting on the Peterson variety $\operatorname{Pet}_{G}$ and we set $H_{S}^{*}(pt)\cong\mathbb{Z}[t]$, where $t$ is a certain character of $S$; see [16, Section 3.2]. By intersecting the Peterson variety $\operatorname{Pet}_{G}$ with Schubert cells and opposite Schubert cells in the Bruhat decomposition (2.1) and opposite Bruhat decomposition (2.2) of the flag variety $G/B$, we have the following
Proposition 2.2 (See [16, Proposition 3.3] and [27, Proposition 6.2]).
@@
For $I\subset\Delta$, set
(2.5) $\operatorname{Pet}_{G,I}:=\overline{\operatorname{Pet}_{G,I}^{\circ}}=\operatorname{Pet}_{G}\cap X_{w_{I}},\qquad\Omega_{I}:=\operatorname{Pet}_{G}\cap\Omega_{w_{I}}.$
$\operatorname{Pet}_{G,I}$ is known as the Peterson Schubert variety in the literature. By the above affine paving (2.4), we have the following theorem; see [32] and [36].

### v2

@@ Abstract.
We give an explicit, positive, and type-uniform formula for all equivariant structure constants of the Peterson Schubert calculus in arbitrary Lie types, using only the Cartan matrix of the corresponding root system $\Phi$. As an application, we derive a type-uniform formula for the mixed $\Phi$-Eulerian numbers.
@@ 1. Introduction
Now, let $G$ be a complex semisimple linear algebraic group with a Borel subgroup $B\subset G$. Recall that the cohomology ring $H^{*}(G/B;\mathbb{Z})$ of the flag variety $G/B$ has a distinguished basis consisting of the Schubert classes [8]. Studying the product structure of these Schubert classes (sometimes also encompassing the study of analogous questions in generalized cohomology theories or equivariant settings) is the central problem in modern Schubert calculus. By Kleiman’s transversality theorem [24], the Schubert structure constants are non-negative integers. In type $A$, the Schubert polynomials in [30] are polynomial representatives of the Schubert classes in the classical Borel presentation [9] of the cohomology ring of the complete flag variety. It is an outstanding problem in combinatorics and enumerative algebraic geometry to give a positive formula or a combinatorial model for the multiplicative structure constants of the Schubert polynomials.
There are also interesting works studying the structure constants of the cohomology ring of the Peterson varieties (the so-called “Peterson Schubert calculus”). Let $\operatorname{Pet}_{G}$ denote the Peterson variety inside $G/B$ (see (2.3) for a precise definition). It admits an action of a one-dimensional sub-torus $S$ of the maximal torus $T\subset G$. Let $\sigma_{v_{I}}\in H_{S}^{*}(G/B)$ be a Schubert class indexed by some Coxeter element $v_{I}\in W$ for $I\subset\Delta$, where $W$ is the Weyl group and $\Delta$ is the set of simple roots. Let $\iota^{*}\left(\sigma_{v_{I}}\right)$ be the equivariant pullback of $\sigma_{v_{I}}$ along the inclusion $\iota:\operatorname{Pet}_{G}\rightarrow G/B$. While the class $\iota^{*}\left(\sigma_{v_{I}}\right)$ depends on the choice of $v_{I}$, the class $p_{I}:=\frac{\iota^{*}\left(\sigma_{v_{I}}\right)}{m\left(v_{I}\right)}$ is independent of this choice after dividing by a certain intersection multiplicity $m\left(v_{I}\right)$, known as the Peterson Schubert class in the literature. In [16], Goldin, Mihalcea, and Singh showed that $\left\{p_{I}\right\}_{I\subset\Delta}$ is an $H_{S}^{*}(pt)$-basis of $H_{S}^{*}(\operatorname{Pet}_{G};\mathbb{Z})$, and that this basis is dual to the equivariant Borel–Moore homology basis constituted by the fundamental classes of the Peterson Schubert cycles (see Theorem 2.5). Furthermore, their main result shows that the structure constants of multiplication $c_{I,J}^{K}\in H_{S}^{*}(pt)\cong\mathbb{Z}[t]$, with respect to the basis $\left\{p_{I}\right\}_{I\subset\Delta}$, defined by
@@
are polynomials in $t$ with non-negative integer coefficients.
A natural open question (asked in [21, P.43] for Lie type $A$) is the following
@@
For Lie type $A$, Goldin–Gorbutt [15] found positive combinatorial formulae for the structure constants in a successive way, while Abe–Horiguchi–Kuwata–Zeng [2] found a different combinatorial model computing these structure constants in non-equivariant cohomology based on a geometric viewpoint. Some years earlier, in type $A$, Harada–Tymoczko [21] gave a positive equivariant Monk formula for the product of a cohomology-degree-2 Peterson Schubert class (that is, $p_{\{i\}}$ for some $i\in\Delta$) with an arbitrary Peterson Schubert class, and Bayegan–Harada [6] gave an equivariant Giambelli formula to express the Peterson Schubert class $p_{I}$ in terms of the ring generators $p_{\{i\}}$’s. Drellich [13] gave the equivariant Monk rule and Giambelli formula for all Lie types.
While the positivity of structure constants in the Peterson Schubert calculus was recently established using the aforementioned geometric and combinatorial methods, the above Question 1.1 remains unresolved in general, and it is not known whether these constants can be calculated directly and uniformly from the intrinsic Lie-theoretic data of the corresponding root system. This paper addresses this problem. In this paper, we give a simple, explicit, positive, and type-uniform formula for all structure constants in (1.1) of the Peterson Schubert calculus in arbitrary Lie types, which completely solves the above question in a type-uniform way. Our formula is a concise algebraic formula, determined only by the Cartan matrix. As far as we know, this is the first example of an explicit, complete, and type-uniform positive algebraic formula in general Schubert calculus. We hope that our method can shed some new light on related structure constants problems in Schubert calculus. We derive this formula from an algebraic viewpoint, which is different from the above-mentioned work. We now present our main results.
Let $C_{\Delta}:=\left(\left\langle\alpha_{i},\alpha_{j}\right\rangle\right)_{i,j\in\Delta}=(c_{ij})_{i,j\in\Delta}$ be the Cartan matrix associated with the semisimple Lie algebra $\mathfrak{g}$ of $G$. Harada–Horiguchi–Masuda [20] gave an explicit and beautiful Borel-type presentation of the equivariant cohomology ring of the Peterson varieties in all Lie types; see Theorem 2.4. The following is our first result.
@@
Using the above Peterson Schubert monomials and the Harada–Horiguchi–Masuda presentation (2.6), we can expand the product of two monomials into the monomial basis representing the Peterson Schubert classes, and compute all the structure constants in (1.1).
To state our main theorem, we first need to introduce some notation. For any $K\subset\Delta$, we use $C_{K}^{-1}]_{j,k}$ to denote the $(j,k)$-entry of the matrix $C_{K}^{-1}$, which is the inverse of the Cartan sub-matrix $C_{K}$ determined by $K$. Suppose $|\Delta|=n$, we use the notation $(\quad)_{J,K}$ to denote a $2^{n}\times 2^{n}$ matrix, with rows indexed by $J\subset\Delta$ and columns indexed by $K\subset\Delta$, and the ordering of the subsets of $\Delta$ is the lexicographical order. The following is our main theorem.
@@
With the notation as above. Suppose $I=\{i_{1},...,i_{l}\}\subset\Delta$. Then the structure constant $c_{I,J}^{K}$, defined in (1.1), is equal to $\frac{\operatorname{det}\left(C_{I}\right)\operatorname{det}\left(C_{J}\right)\left|W_{K}\right|}{\left|W_{I}\right|\left|W_{J}\right|\operatorname{det}\left(C_{K}\right)}$ multiplied by the $(J,K)$-entry of the product of matrices $(d^{K}_{i_{1},J})_{J,K}\cdots(d^{K}_{i_{l},J})_{J,K}$, where
@@
Since the ordinary cohomology $H^{*}(\operatorname{Pet}_{G};\mathbb{Z})$ vanishes in odd degrees, the Peterson variety $\operatorname{Pet}_{G}$ is equivariantly formal in the sense of [18]. Therefore, by letting $t=0$ in (1.3), we get a simple formula for all non-equivariant structure constants with respect to the basis $\left\{p_{I}\right\}_{I\subset\Delta}$ of $H^{*}(\operatorname{Pet}_{G};\mathbb{Z})$ (we use the same notation to denote the classes under the natural map $H_{S}^{*}(\operatorname{Pet}_{G};\mathbb{Z})\rightarrow H^{*}(\operatorname{Pet}_{G};\mathbb{Z})$).
@@
Suppose $I=\{i_{1},...,i_{l}\}\subset\Delta$. Then the non-equivariant structure constant $n_{I,J}^{K}$, defined to be the constant term of $c_{I,J}^{K}$ in (1.1), is equal to $\frac{\operatorname{det}\left(C_{I}\right)\operatorname{det}\left(C_{J}\right)\left|W_{K}\right|}{\left|W_{I}\right|\left|W_{J}\right|\operatorname{det}\left(C_{K}\right)}$ multiplied by the $(J,K)$-entry of the product of matrices
$(m^{K}_{i_{1},J})_{J,K}\cdots(m^{K}_{i_{l},J})_{J,K}$, where
@@
To prove the above theorems, we transform the structure constants in (1.3) into an infinite matrix series (see the proof of Theorem 3.5) related to the Coxeter adjacency matrix $A_{K}$, defined as $A_{K}:=2E-C_{K}$, where $E$ is the identity matrix. We use the classical Perron–Frobenius theorem to prove that this matrix series is convergent and related to the inverse of a matrix; see Lemma 3.7. Finally, we use the Sherman–Morrison formula to compute the matrix inverse, which is simplified to the final formula; see Lemma 3.8.
It is well known that the entries of the inverse of any Cartan matrix are all non-negative (see, for example, [23, Exercise 13.8]), hence the entries of all the matrices $(d^{K}_{i,J})_{J,K}$ in Theorem 1.3 are polynomials in $t$ with non-negative coefficients. Since the entries in the product of such matrices remain polynomials in $t$ with non-negative coefficients, our formula is indeed a positive formula, which provides a new purely algebraic proof of the positivity of equivariant structure constants of the Peterson Schubert calculus. We also provide a simple criterion for when these structure constants are non-zero; see Corollary 3.13 and Remark 3.14.
The relationship between the (non-equivariant) Peterson Schubert calculus and the mixed Eulerian numbers (which were introduced and studied by Postnikov [34]) has been actively explored very recently; see [7], [32], and [22]. These numbers compute the mixed volumes of $\Phi$-hypersimplices, see Section 4 for the precise definitions. The origins of this connection lie in the realization of the Peterson variety as a flat degeneration of a smooth projective toric variety, the $W$-permutohedral variety; see, for example, [1, 25, 26]. In [22], Horiguchi provided a combinatorial model introduced in [2] for the computation of the mixed $\Phi$-Eulerian numbers when $\Phi$ is of type $A$ and derived a type-by-type computation for the mixed $\Phi$-Eulerian numbers for general Lie types by iteratively applying the (non-equivariant version of) Monk formula from Drellich in [13]. As an application of our main theorem, we derive a type-uniform formula for the mixed $\Phi$-Eulerian numbers in arbitrary Lie type.
@@
Let $\Phi$ be an irreducible root system of rank $n$. Let $c_{1},\ldots,c_{n}$ be non-negative integers with $c_{1}+\cdots+c_{n}=n$. The mixed $\Phi$}-Eulerian number $A_{c_{1},\dots,c_{n}}^{\Phi}$ can be computed using the following formula:
@@
Compared with the computations in [22], our formula for the mixed $\Phi$-Eulerian numbers is quite simple and direct, which avoids the need to discuss the changes in the Lie types case-by-case when considering sub-root systems as in [22, Section 7].
The remainder of the paper is organized as follows. In Section 2, we introduce the notation and some preliminary results which we use throughout this paper. In Section 3, we prove our main theorems. In Section 4, we give the application to the mixed Eulerian numbers.
@@ Acknowledgments
The first author would like to express his gratitude to Hiraku Abe and Haozhi Zeng for helpful discussions. The authors thank Yibo Gao for some useful comments. The first author is partially supported by NSFC Grant No. 12471309.
@@ 2. Preliminaries
In this section, we collect our notation and preliminary results for later use.
@@ 2.1. Flag varieties and Schubert varieties
Let $G$ be a complex connected, simply connected, semisimple algebraic group of rank $n$, $B\subset G$ be a Borel subgroup, and $B^{-}\subset G$ be the opposite Borel. Write $T:=B\cap B^{-}$ for the maximal torus common to both Borel subgroups. Let $\mathfrak{g}=\operatorname{Lie}(G)$, $\mathfrak{b}=\operatorname{Lie}(B)$, and $\mathfrak{h}=\operatorname{Lie}(T)$ be the corresponding Lie algebras. Denote by $\Phi$ the root system determined by $(G,T)$, and by $\Delta=\{\alpha_{1},\dots,\alpha_{n}\}$ the simple roots specified by $(B,T)$. We abuse the notation by letting $\Delta$ also denote the index set of simple roots. The Weyl group $W:=\langle s_{i}\mid i\in\Delta\rangle$ is a finite Coxeter group. For any subset $I\subset\Delta$, we have the standard parabolic subgroup $P_{I}\subset G$, whose Weyl group is $W_{I}=\langle s_{i}\mid i\in I\rangle$, a parabolic subgroup of $W$.
@@
the Schubert and opposite Schubert varieties in $G/B$, respectively. It follows from (2.2) that the equivariant cohomology classes $\sigma_{w}\in H_{T}^{*}(G/B;\mathbb{Z})$, which are Poincaré dual to the fundamental class $\left[\Omega_{w}\right]$, form a basis $\left\{\sigma_{w}\mid w\in W\right\}$ of $H_{T}^{*}(G/B;\mathbb{Z})$ as a module over $H_{T}^{*}(pt)$, called the equivariant Schubert classes.
@@ 2.2. Peterson varieties and some geometric constructions
(2.3) $\operatorname{Pet}_{G}:=\Bigl\{gB\in G/B\;\Big|\;\text{Ad}_{g^{-1}}(e)\in\mathfrak{b}\oplus{\textstyle\bigoplus\limits_{\alpha\in\Delta}}\mathfrak{g}_{-\alpha}\Bigr\}.$
It is known that the Peterson variety $\operatorname{Pet}_{G}$ is always irreducible with complex dimension $n:=\operatorname{rank}G$ (see [33] or [28, Proposition 6.2]), but it is singular and even non-normal in general [27]. The Peterson varieties also form a special class of a larger family of subvarieties of the flag variety, namely the Hessenberg varieties introduced in [12]; see [3] for a very nice survey.
@@
where $NV_{i}$ denotes the image of $V_{i}$ under the linear map $N\colon\mathbb{C}^{n}\to\mathbb{C}^{n}$. In this setting, the Peterson varieties are also known as the regular nilpotent Hessenberg varieties with Hessenberg function $h(i)=i+1$ for $i=1,2,\ldots,n-1$.
It is known that there is a one-dimensional sub-torus $S\subset T$ acting on the Peterson variety $\operatorname{Pet}_{G}$ and we set $H_{S}^{*}(pt)\cong\mathbb{Z}[t]$, where $t$ is a certain character of $S$; see [16, Section 3.2]. By intersecting the Peterson variety $\operatorname{Pet}_{G}$ with Schubert cells and opposite Schubert cells in the Bruhat decomposition (2.1) and opposite Bruhat decomposition (2.2) of the flag variety $G/B$, we have the following
Proposition 2.2 (See [16, Proposition 3.3] and [28, Proposition 6.2]).
@@
For $I\subset\Delta$, we set
(2.5) $\operatorname{Pet}_{G,I}:=\overline{\operatorname{Pet}_{G,I}^{\circ}}=\operatorname{Pet}_{G}\cap X_{w_{I}},\qquad\Omega_{I}:=\operatorname{Pet}_{G}\cap\Omega_{w_{I}},$
where $\operatorname{Pet}_{G,I}$ is known as the Peterson Schubert variety in the literature. By the above affine paving (2.4), we have the following theorem; see [33] and [37].